OpenStudy (anonymous):

Suppose you know the graph of f contains the point (2,3). If g(x)=f(x+5), what point must be on the graph of g? Possible answers: (7,2), (2,-2), (-3,3), (3,2), (2,8)

OpenStudy (anonymous):

Is it (2,8)?

OpenStudy (anonymous):

when x + 5 = 2 you know x = -3 right?

OpenStudy (anonymous):

in other words solve \[x+5=2\] gives \[x=-3\] so \[g(-3)=f(-3+5)=f(2)=3\] telling you that \[(-3,3)\] is on the graph of g

OpenStudy (anonymous):

Oooh okay thanks!

OpenStudy (anonymous):

yw